OpenStudy (anonymous):

Find the quotient using the rules for exponents: 86 ÷ 82

OpenStudy (anonymous):

wait its actually 8^6 ÷ 8^2

OpenStudy (across):

I was about to say. ^^

OpenStudy (across):

Here's a hint. It's pretty self-explanatory:\[\frac{a^b}{a^c}=a^{b-c}\]:)

OpenStudy (anonymous):

do i subtract on this one too? 19^3 × 19^4

OpenStudy (across):

On that one, the rule is:\[a^b\cdot a^c=a^{b+c}\]:)

OpenStudy (anonymous):

so i add for multiplication and i subtract for division? and is there a plus and minus one too? if so whats the formula for them?

OpenStudy (across):

That's correct, and unfortunately, there are no exponent rules for addition and subtraction of powers. That is,\[a^b+a^c\text{ and}\\a^b-a^c\]is as good as you can get them to be.

OpenStudy (across):

are* gosh, my English slips at times!

OpenStudy (anonymous):

Oh ok thank you very much now i understand hha! :)

OpenStudy (across):

You're welcome. :)